geerky42 (geerky42):

The sides of a right triangle form a primitive Pythagorean triple. If its semi perimeter is 120, find its area.

OpenStudy (tkhunny):

So, A + B + C = 240 = 2*S \(A^{2} + B^{2} = C^{2}\) \(Area = \sqrt{S\cdot(S-A)\cdot(S-B)\cdot(S-C)}\) Anything looking familiar?

geerky42 (geerky42):

What about AB/2 = Area?

OpenStudy (tkhunny):

So much information! How could you NOT find a solution?!

geerky42 (geerky42):

I know this looks easy and I got 2400, but correct answer is 840... I just don't understand...

geerky42 (geerky42):

I can see that A= 80, B= 60 and C = 100 I don't ever know where 840 comes from...